The answer and equation to how many liters of 80% acid solution must be mixed with a 20% acid solution to get 600 L of a 70% as a solution
Answer by fractalier(6550)

Let x be the amount of 80% acid solution.
Thus 600-x will be the amount of 20% acid.
The set up is then
.80x + .20(600-x) = .70(600)
.80x + 120 - .20x = 420
.6x = 300
x = 500 L of 80% solution
600-x = 100 L of 20% solution